# Hydrogen-Alpha as a Tracer of Star Formation in the SPHINX Cosmological Simulations

### v1

@@ 3.2 Simulation-informed SFR() calibrations
Using the insights gained from our analysis of the SFR— relation in SPHINX, we discuss possible improvements to the existing SFR() calibrations that take the form of Eq. 1. For example, stellar metallicity could be a useful parameter to include in the calibration to reduce the error in the predicted SFR (see Fig. 2). However, measuring stellar metallicity is difficult in practice, therefore we only use quantities that can be measured robustly from galaxy spectra. Specifically, we run an ordinary least squares (OLS) model similar to Eq. 1 but allowing the coefficient to vary. We use the randomly selected of the sample to train the model and the remaining to validate its performance. The resulting SFR() calibration can be written as follows:
where the last term is the effective correction to Eq. 1. We find that the root mean squared error (RMSE) calculated on the training and test datasets is identical ( dex), indicating that the model performs well on unseen data. Compared to the T19 calibration () which yields dex calculated on the full dataset, the RMSE is reduced by dex (see also Fig. 3).
The SFR() calibration given in Eq. 2 corrects the SFRs by a factor that depends on only, which means that is neglected even when using the updated calibration. To further improve on this, we add a correction to Eq. 2 that depends on the equivalent width (). This choice is motivated by the fact that the traces stellar metallicity and age that both drive (Fig. 2). In addition, the is only weakly sensitive to dust attenuation, which implies that dust corrections have a smaller impact on the results compared to when other observables are used (e.g., ). We therefore run an OLS model that includes , using the same training sample as in Eq. 2. The resulting SFR() calibration can be written as follows:
where the last two terms are the effective corrections to Eq. 1. The new calibration yields dex on both the training and test datasets, dex lower than the RMSE calculated using Eq. 2, or dex lower than the RMSE calculated using the T19 calibration (). Figure 3 compares these calibrations as a function of . In particular, this plot shows that the median in the bins is within dex across the entire range of luminosities () when using Eq. 2 or Eq. 3. This is an improvement over the T19 calibration, which overpredicts the SFRs by dex in galaxies with low-to-medium luminosities (). The standard deviation of in the bins (hereafter ) is nearly identical between the T19 calibration and the SFR() calibration given in Eq. 2 ( dex), but it decreases by dex ( dex) when using Eq. 3, especially at medium luminosities ().
@@ 4.1 Implications
The line has long been known as a reliable indicator of the SFR as it traces the ionizing photon luminosity from massive (), short-lived ( Myr) stars (e.g., Kennicutt, 1998). Over the past decades, has been widely employed in numerous spectroscopic (e.g., Shim et al., 2009; Gunawardhana et al., 2013; Nagaraj et al., 2023) and narrow-band imaging (e.g., Ly et al., 2007; Hayes et al., 2010; Sobral et al., 2013) surveys to measure the cosmic star formation history (SFH) up to . With the recent advent of sensitive near-IR spectroscopy at – m, these measurements have been extended to –, in particular thanks to the JWST/NIRCam slitless spectroscopic observations of large samples of -emitters at the same redshifts (e.g., Covelo-Paz et al., 2025; Fu et al., 2025; Lin et al., 2025).
@@
As observations become more accessible at high redshifts, it is crucial to revisit the classical SFR() calibrations (e.g., Kennicutt, 1998) commonly used at . Based on a realistic galaxy simulation, we have shown that such calibrations bias the predicted SFRs under the physical conditions that are characteristic of high-redshift galaxies observed with JWST (i.e., low metallicity and bursty star formation), potentially also affecting population statistics. To quantify this effect, we assess the impact of the metallicity- and age-dependent SFR bias (see Fig. 2) on the cosmic star formation rate density () measurements. Specifically, we convert the dust-corrected luminosity function at presented by Fu et al. (2025) to the star-formation rate function (SFRF) using the T19 calibration () and our new SFR() calibration given in Eq. 2 which, by design, takes into account the evolution of the SFR— relation with metallicity. Following previous studies, we integrate both SFRFs down to yr-1 which corresponds to (Bouwens et al., 2015). We find that the resulting decreases from to , or by %, when using the new calibration. This suggests that the influence of the evolving SFR— relation on is moderate, likely because galaxies where the SFR bias is highest contribute little to . For example, galaxies with contribute less than % to the total . If we integrate the SFRF only up to , then the change in would be more pronounced: would decrease by %.
@@
The SFR bias also impacts the correlations between the SFR and other galaxy properties, such as stellar mass. In Fig. 4, we show the SFR— relation, also known as the galaxy main sequence (Noeske et al., 2007), in the JWST All the Little Things (ALT) survey (Naidu et al. 2024, Di Cesare et al. in prep.). We find that the slope of the SFMS increases by (%) when we use Eq. 3 instead of the T19 calibration to calculate the SFRs. Equation 2 yields similar results, although with a smaller significance (). In addition, we find that the scatter in the SFMS decreases by dex (%) when using Eq. 3, most noticeably at the low-mass end ( dex at ). In contrast, Eq. 2 produces almost the same scatter as the T19 calibration ( dex). We note that the SMFS in Fig. 4 is impacted by selection effects: some low-mass galaxies are not detected in observations due to their faintness, resulting in a shallow SMFS slope. However, proper modeling of these selection biases is out of scope of this paper and will be discussed elsewhere (see Di Cesare et al. in prep.).
@@ 4.2 Dust attenuation
In Sect. 3, we ignored the effects of dust attenuation, assuming that the H line measurements (i.e., and ) can be corrected for such effects without errors. This assumption is also usually made for the SFR() calibrations presented in the literature. However, dust attenuation is known to be an important source of systematic error in the -derived SFRs (e.g., Hopkins et al., 2001). This effect likely persists even at as the amount of dust in -emitters at these redshifts is found to be non-negligible (e.g., Covelo-Paz et al., 2025 reported the median mag for a sample of -emitters at ).
Here, we test the possibility of deriving SFRs directly from the observed emission without prior dust corrections. We use the dust-attenuated luminosity () and equivalent width () from SPHINX, calculated as described in Sect. 2, and fit the same linear relations between the SFR, , and as in Sect. 3. We present the resulting SFR() calibrations in App. B. The calibration that depends on only (Eq. 4; Fig. 5, cyan) yields dex, while the calibration that additionally depends on (Eq. 5; Fig. 5, magenta) yields dex. In both cases, the RMSE is significantly higher compared to the RMSEs calculated using the calibrations that are based on the intrinsic properties (see Sect. 3). We also find that the SFRs are significantly underestimated at high luminosities (), with Eq. 4 and Eq. 5 producing the median and , respectively. This can be explained by the fact that high- galaxies in SPHINX have, on average, very high attenuation values () which we are unable to fully correct for while simultaneously trying to model the SFRs of low- galaxies. However, we caution that such are not reported in high-redshift observations (see Sect. 2) and thus making accurate SFR estimates based on the dust-attenuated might still be possible in practice.
We also test whether the Balmer decrement (; hereafter ), whose measurements are commonly used to calculate the dust corrections for , can be a useful parameter to include in the SFR() calibration. Interestingly, we find that the -based calibration (Eq. 6; Fig. 5, orange) yields dex, which is dex higher than the RMSE calculated using the -based calibration. This indicates that is not as powerful as in constraining the SFRs of SPHINX galaxies. Including both and in the calibration (Eq. 7; see also Fig. 5, light blue) yields the best results ( dex). Observationally, however, the measurements are hampered by slitlosses and the requirement of a broader wavelength coverage. This implies that the benefits of including in the SFR() calibration are likely limited, with remaining a key addition to the calibration.
@@ 4.3 Caveats
Extending the calibrations to erg s-1 is already possible with SPHINX but a larger halo catalog compared to the currently available SPDRv1 catalog is needed to create a complete sample down to very low luminosities. At low redshifts (), the SFHs become less bursty and metallicities increase, rendering the classical calibrations more suitable. Finally, at high redshifts (), it remains to be established whether the new calibrations can be used, as the physical conditions of star formation at such redshifts are poorly understood.
We also stress the fact that the new calibrations assume the BPASSv SPS model with a Kroupa-type IMF (see Sect. 2). Figure 1 shows that the choice of the IMF significantly impacts the predicted SFRs, resulting in a systematic error of up to dex. The choice of the SPS model, in particular whether the model includes binary stars or not, results in a smaller systematic error ( dex), possibly because binary stars have little effect on the LyC luminosity in the first 3 Myrs of a stellar population (e.g., Rosdahl et al., 2018), however such errors are still non-negligible. In the future, these systematic errors can be reduced by testing the SPS models and constraining the IMF at , e.g., by means of deep rest-frame UV spectroscopy (e.g., Steidel et al., 2016; Chisholm et al., 2019).
Finally, we caution that SPHINX does not include black hole formation due to its small volume. This implies that the AGN contribution to the emission is neglected in the new SFR() calibrations. Hence, it is necessary to remove the AGN from the sample before applying these calibrations.
@@ 5 Summary
In this work, we study the relationship between emission and the SFR in the SPHINX cosmological simulations at . Using the simulated galaxy data from the SPHINX Public Data Release, Version 1 (Katz et al., 2023), we select a sample of star-forming galaxies that are representative of the -emitter population routinely observed with JWST at . We find that the median SFR— relation in SPHINX exhibits a significant evolution (Fig. 1), with a downturn at low luminosities (–) that is consistent with fainter galaxies having low stellar metallicites () and thus being more efficient producers of ionizing photons than brighter, relatively metal-enriched () galaxies.
Due to the metallicity dependence of the SFR— relation, the classical SFR() calibrations (e.g., Kennicutt, 1998; Eq. 1), on average, overestimate the SFRs of faint galaxies in SPHINX by dex (Fig. 3, black). In Eq. 2, we propose a new SFR() calibration that depends on the intrinsic luminosity only and reduces the error in the predicted SFRs by dex (Fig. 3, cyan). In Eq. 3, we propose another SFR() calibration that additionally depends on the equivalent width, a parameter that is sensitive to the stellar metallicity and age (Fig. 2). This calibration further improves the SFR estimate, reducing the error by a total of dex (Fig. 3, magenta). The new SFR() calibrations affect the galaxy population statistics at high redshifts (): the inferred cosmic star-formation density descreases by %, and the slope of the star-formation main sequence increases by (Fig. 4).

### v2

@@ 3.2 Simulation-informed SFR() calibrations
Using the insights gained from our analysis of the SFR— relation in SPHINX, we discuss possible improvements to the existing SFR() calibrations that take the form of Eq. 1. For example, stellar metallicity could be a useful parameter to include in the calibration to reduce the error in the predicted SFR (see Fig. 2). However, measuring stellar metallicity is difficult in practice, and we therefore only use quantities that can be measured robustly from galaxy spectra. Specifically, we run an ordinary least-squares (OLS) model similar to Eq. 1, but allowing the coefficient to vary. We use the randomly selected of the sample to train the model and the remaining to validate its performance. The resulting SFR() calibration can be written as follows:
where the last term is the effective correction to Eq. 1. The root mean squared error (RMSE) calculated on the training and test datasets is identical ( dex), indicating that the model performs well on unseen data. Compared to the T19 calibration () which yields dex calculated on the full dataset, the RMSE is reduced by dex (see also Fig. 3).
The SFR() calibration given in Eq. 2 corrects the SFRs by a factor that depends on alone, which means that is neglected even when using the updated calibration. To further improve on this, we add a correction to Eq. 2 that depends on the equivalent width (). This choice is motivated by the fact that traces stellar metallicity and age, which both drive (Fig. 2). In addition, is only weakly sensitive to dust attenuation, which implies that dust corrections affect the results less than when other observables are used (e.g., ). We therefore run an OLS model that includes , using the same training sample as in Eq. 2. The resulting SFR() calibration can be written as follows:
where the last two terms are the effective corrections to Eq. 1. The new calibration yields dex on the training and test datasets, which is dex lower than the RMSE calculated using Eq. 2, or dex lower than the RMSE calculated using the T19 calibration (). Figure 3 compares these calibrations as a function of . In particular, this plot shows that the median in the bins is within dex across the entire range of luminosities () when using Eq. 2 or Eq. 3. This is an improvement over the T19 calibration, which overpredicts the SFRs by dex in galaxies with low to medium luminosities (). The standard deviation of in the bins (hereafter ) is nearly identical between the T19 calibration and the SFR() calibration given in Eq. 2 ( dex), but it decreases by dex ( dex) for Eq. 3, especially at medium luminosities ().
We also run a similar model in which we replace with , which is a proxy for the gas-phase (and, to a weaker extent, stellar) metallicity. The resulting calibration shows the same performance as Eq. 3 ( dex). Given the broader wavelength coverage required to measure compared to , we find Eq. 3 to be more practical when applied to observations and hence use it as our best-performance calibration for the remainder of this paper.
@@ 4.1 Implications
The line has long been known as a reliable indicator of the SFR as it traces the ionizing photon luminosity from massive (), short-lived ( Myr) stars (e.g., Kennicutt, 1998). Over the past decades, has been widely employed in numerous spectroscopic (e.g., Shim et al., 2009; Gunawardhana et al., 2013; Nagaraj et al., 2023) and narrow-band imaging (e.g., Ly et al., 2007; Hayes et al., 2010; Sobral et al., 2013) surveys to measure the cosmic SFH up to . With the recent advent of sensitive near-IR spectroscopy at – m, these measurements have been extended to –, in particular thanks to the JWST/NIRCam slitless spectroscopic observations of large samples of -emitters at the same redshifts (e.g., Covelo-Paz et al., 2025; Fu et al., 2025; Lin et al., 2026).
@@
As observations become more accessible at high redshift, it is crucial to revisit the classical SFR() calibrations (e.g., Kennicutt, 1998) that are commonly used at . Based on a realistic galaxy simulation, we have shown that these calibrations bias the predicted SFRs under the physical conditions that are characteristic of high-redshift galaxies observed with JWST (i.e., a low metallicity and bursty star formation), potentially also affecting population statistics. To quantify this effect, we study the effect of the metallicity- and age-dependent SFR bias (see Fig. 2) on the measurements. Specifically, we convert the dust-corrected luminosity function at presented in Schechter form by Fu et al. (2025) into the star formation rate function (SFRF). We perform this conversion by expressing the luminosity via SFR using two SFR— calibrations: the T19 calibration (), and our new SFR() calibration given in Eq. 2, which, by design, takes the dependence of the SFR— relation on metallicity into account. Following previous studies, we then integrate both SFRFs down to yr-1 which corresponds to (Bouwens et al., 2015). The resulting decreases from to , or by %, when using the new calibration. This suggests that the effect of the evolving SFR— relation on is moderate, likely because galaxies in which the SFR bias is highest contribute little to . For example, galaxies with contribute less than % to the total . If we integrate the SFRF only up to , the change in would be more pronounced: would decrease by %.
@@
The SFR bias also affects the correlations between the SFR and other galaxy properties, such as stellar mass. In Fig. 4, we show the SFR— relation, also known as the star formation main sequence (SFMS; Noeske et al., 2007), in the JWST All the Little Things (ALT) survey (Naidu et al., 2024; Di Cesare et al., 2025). We apply different SFR() calibrations to the dust-corrected luminosities and calculate the slope of the SFMS. The slope increases by (%) when we use Eq. 3 instead of the T19 calibration to calculate the SFRs. Equation 2 yields similar results, although they are less significant (). In addition, we find that the scatter in the SFMS decreases by dex (%) when using Eq. 3, most noticeably at the low-mass end ( dex at ). In contrast, Eq. 2 produces almost the same scatter as the T19 calibration ( dex). We note that the SFMS in Fig. 4 is affected by selection effects: some low-mass galaxies are not detected in observations due to their faintness, resulting in a shallow SFMS slope. Temporary cessation or strong suppression of star formation in galaxies during so-called mini-quenching events (e.g., Gelli et al., 2023; Looser et al., 2025) can further amplify this bias. However, a comprehensive discussion of caveats associated with the SFMS slope measurements is beyond the scope of this paper (this discussion can be found in Di Cesare et al., 2025).
@@ 4.2 Dust attenuation
In Sect. 3, we ignored the effects of dust attenuation, assuming that the H line measurements (i.e., and ) can be corrected for these effects without errors. This assumption is also usually made for the SFR() calibrations presented in the literature. However, dust attenuation is known to be an important source of systematic error in the -derived SFRs (e.g., Hopkins et al., 2001). This effect likely persists even at as the number of dust in -emitters at these redshifts is found to be non-negligible (e.g., Covelo-Paz et al., 2025 reported the median mag for a sample of -emitters at ).
Here, we test the possibility of deriving SFRs directly from the observed emission without prior dust corrections. We use the dust-attenuated luminosity () and equivalent width () from SPHINX, calculated as described in Sect. 2, and fit the same linear relations between the SFR, , and as in Sect. 3. We present the resulting SFR() calibrations in App. B. The calibration that only depends on (Eq. 4; Fig. 5, cyan) yields dex, while the calibration that additionally depends on (Eq. 5; Fig. 5, magenta) yields dex. In both cases, the RMSE is significantly higher than the RMSEs calculated using the calibrations that are based on the intrinsic properties (see Sect. 3). Moreover, the SFRs are significantly underestimated at high luminosities (), with Eq. 4 and Eq. 5 producing the median and , respectively. This can be explained by the fact that high- galaxies in SPHINX have very high attenuation values () on average, which we are unable to fully correct for while simultaneously trying to model the SFRs of low- galaxies. However, we caution that like this are not reported in high-redshift observations (see Sect. 2), and it might therefore still be possible in practice to make accurate SFR estimates based on the dust-attenuated .
We also test whether the Balmer decrement (; hereafter ), whose measurements are commonly used to calculate the dust corrections for , can be a useful parameter to include in the SFR() calibration. Interestingly, we find that the -based calibration (Eq. 6; Fig. 5, orange) yields dex, which is dex higher than the RMSE calculated using the -based calibration. This indicates that is not as powerful as in constraining the SFRs of the SPHINX galaxies. This result is surprising because is commonly regarded as a robust indicator of dust attenuation (), while is in principle mostly independent of it. However, we still find a negative correlation between and in SPHINX (, ; Fig. 6), possibly because the amount of dust increases with the age of the stellar population, a parameter to which is sensitive. Further, unlike , correlates with stellar metallicity, allowing for an even more accurate prediction of the SFRs.
Inclusion of both and in the calibration (Eq. 7; see also Fig. 5, light blue) provides the best performance among the considered models ( dex). Observationally, however, the measurements are hampered by slit losses and the requirement of a broader wavelength coverage. This implies that the benefits of including in the SFR() calibration are likely limited, with remaining a key addition to the calibration.
@@ 4.3 Caveats
It is already possible to extend the calibrations to erg s-1 with SPHINX, but a larger halo catalog than the currently available SPDRv1 catalog is needed to create a complete sample down to very low luminosities. At low redshifts (), the SFHs become less bursty and the metallicities increase, rendering the classical calibrations more suitable. Finally, at high redshift (), it remains to be established whether the new calibrations can be used, as the physical conditions of star formation at these redshifts are poorly understood.
We also stress the fact that the new calibrations assume the BPASSv SPS model with a Kroupa-type IMF (see Sect. 2). Figure 1 shows that the choice of the IMF significantly affects the predicted SFRs, resulting in a systematic error of up to dex (see also Jeřábková et al., 2018, who studied the effect of a non-universal IMF on the SFR— relation). The choice of the SPS model, in particular, whether the model includes binary stars, results in a smaller error ( dex), possibly because binary stars have little effect on the LyC luminosity in the first Myrs of a stellar population when the bulk of ionizing photons is produced (e.g., Rosdahl et al., 2018), but this effect is still non-negligible. In the future, these systematic errors can be reduced by testing the SPS models and constraining the IMF at , for instance, by means of deep rest-frame UV spectroscopy (e.g., Steidel et al., 2016; Chisholm et al., 2019).
@@ 5 Summary
In this work, we study the relation between emission and the SFR in the SPHINX cosmological simulations at . Using the simulated galaxy data from the SPHINX Public Data Release, Version 1 (Katz et al., 2023), we select a sample of star-forming galaxies that are representative of the -emitter population routinely observed with JWST at . The median SFR— relation in SPHINX exhibits a downturn at low luminosities (Fig. 1; –). This downturn is consistent with fainter galaxies having lower stellar metallicities () and thus being more efficient producers of ionizing photons than brighter, relatively metal-enriched () galaxies. Furthermore, the SFR— relation exhibits a significant scatter at fixed luminosity, ranging from dex at to dex at (i.e., the scatter is larger at lower luminosities). We demonstrate that this scatter is primarily driven by the variations in the stellar metallicities and ages (Fig. 2), whereas other properties (e.g., and ) play a less significant role.
Because of the metallicity dependence of the SFR— relation, the classical SFR() calibrations (e.g., Kennicutt, 1998; Eq. 1) overestimate the SFRs of faint galaxies in SPHINX on average by dex (Fig. 3, black). In Eq. 2, we propose a new SFR() calibration that depends on the intrinsic luminosity alone and reduces the error in the predicted SFRs by dex (Fig. 3, cyan). To address the scatter in the SFR— relation, we propose another SFR() calibration in Eq. 3. This calibration additionally depends on the equivalent width, a parameter that is sensitive to the stellar metallicity and age (Fig. 2), and further improves the SFR estimate, reducing the error by a total of dex (Fig. 3, magenta). The new SFR() calibrations affect the galaxy population statistics at high redshift (): the inferred decreases by %, and the slope of the SFMS increases by (Fig. 4).